OpenStudy (anonymous):

The work of a student trying to solve the equation 2(2x − 3) = 6 + 2x + 4 is shown below: Step 1: 2(2x − 3) = 6 + 2x + 4 Step 2: 4x − 3 = 10 + 2x Step 3: 4x − 2x = 10 + 3 Step 4: 2x = 13 Step 5: x = 6.5 In which step did the student first make an error and what is the correct step? Step 2: 4x − 6 = 10 + 2x Step 2: 4x − 3 = 2 + 2x Step 3: 4x + 2x = 10 + 3 Step 3: 4x − 2x = 10 - 3

OpenStudy (anonymous):

i got the first choice for my answer

OpenStudy (calculusxy):

Yes :)

OpenStudy (anonymous):

are u sure ?

OpenStudy (calculusxy):

Yeah because 2(2x - 3) = 2(2x) - 2(3) = 4x - 6

OpenStudy (anonymous):

ok thanks :D

OpenStudy (calculusxy):

You're welcome :)